In a recent​ year, the scores for the reading portion of a test were normally​ distributed, with a mean of 22.6 and a standard deviation of 6.8 Complete parts​ (a) through​ (d) below.

b. find the probability that a randomly selected high school student who took the reading portion of the test has a score that is between 17.3 and 27.9
Answer #1

Let mean = 22.6 and standard deviation o = 6.8 for normal distribution

The probability that a randomly selected high school student test

has a score between 17.3 and 27.9

  Let I1 = 17.3 and 22 = 27.9

  0-2823 : 0.2823 Z=-0478 220-78

p(21-<<< *=4) = P(17.3- 226 <= < 27.9=226) T 6.8 6.8

  = P(-0.78 < 2 <0.78)

   = Area from 2=0 to z = -0.78 + area from 2=0 to z = 0.78

= 0.2823 +0.2823

  = 0.5646